Lab Notebook Trial 1 Trial 2 Trial 3 100 ml 100 ml 23 °C 100 mL 24 C 38 c 278 mol 25 mol Volume of HCI Initial temperature of HCI Volume of NaOH Initial temperature of NaOH FInal solution temperature Moles of H Moles of OH Limiting reagent Heat transfer for solution 100 mL 24 C 100 mL 24 C 40℃ 278 mol 25 mol 22℃ 100 ml 23 C 38℃ 278 mol 25 mol OH OH OH HC kJ kJ Enthalpy of neutralization per mole of limiting reagent k/mol K/mol Mean enthalov of neutralization per mole of limiting reagent Redo Take
media%2Fdac%2Fdac6725e-c2a3-4504-80eb-b1
media%2F223%2F2233b1e2-be33-4826-8da1-21
Answer #1

The specific heat of water is 4.184 J/g0C. The density of both HCL and NaOH is 1.00 g/mL

Assume that specific heats HCL and NaOH are also 4.1984 J/g C.

For Trial 1

To calculate the heat actually evolved, use the formula

                                            q = m c Δt

To convert ml to grams, multiply volume with density (m = V X D)

(100 mL + 100 mL) = 200 mL of solution

100 mL X 1 g/ml = 2 00 grams of solution.

Find the temperature change.

Δt = t final – t initial = 40 0C – 24 0C =160C

q = m c Δt

q = 200 grams X 4.184   J /g0C x 16 0C

q= 13388.8 J

q = 13.39KJ

This is the heat gained by the water, but in fact it is the heat lost by the reacting HCl and NaOH, therefore q = - 13.39KJ    

Moles of HCl = 0.278

Moles of NaOH = 0.25   

NaOH is limiting reagent   

Enthalpy of neutralization /mol of limiting reagent = -13.39KJ /0.25     

   =- 52.56 KJ/mol  

For Trial 2

Δt = t final – t initial = 38 0C – 23.5 0C =14.50C

q = m c Δt

q = 200 grams X 4.184   J /g0C x 14.5 0C

q= -12133.6 J

q = - 12.13KJ      

Moles of HCl = 0.278

Moles of NaOH = 0.25                 

Enthalpy of neutralization /mol of limiting reagent = -12.13KJ /0.25     

     =- 48.52 KJ/mol  

For Trial 3

Δt = t final – t initial = 38 0C – 22.5 0C =15.5 0C

q = m c Δt

q = 200 grams X 4.184   J /g0C x 15.5 0C

q= 12970.4 J

q = -12.97KJ   

Enthalpy of neutralization /mol of limiting reagent =- 12.97KJ /0.25     

     = -51.88 KJ/mol  

Mean enthalpy of neutralization /mol of limiting reagent

= (-52.56 KJ/mol   + -48.52 KJ/mol   +-51.88 KJ/mol ) /3

Mean enthalpy of neutralization /mol of limiting reagent - 50.99 KJ/mol